OpenStudy (anonymous):

simplify: 4x/9 + x-10/9 ?

OpenStudy (anonymous):

\[\frac{4x}{9} + x - \frac{10}{9}\]Multiply the middle x term by 9/9 to get a common denominator. Then you can combine the numerators and write the expression and a single fraction.

jhonyy9 (jhonyy9):

(4x +9x-10)/9 = (13x -10)/9